OpenStudy (anonymous):

Evaluate: -6 - (-4) -10 -2 24 2

OpenStudy (akashdeepdeb):

negaive times negative is POSITIVE! -6-(-4) = -6 + 4 = -2 [Because -6 is greater in number with a negative sign]

OpenStudy (akashdeepdeb):

Understood @hicks ?

OpenStudy (anonymous):

yes its d

OpenStudy (akashdeepdeb):

B) -2 :)

OpenStudy (anonymous):

oh im confused and sad :)

OpenStudy (anonymous):

thankyou very much

OpenStudy (akashdeepdeb):

:) Be happy.

OpenStudy (anonymous):

im happy :)

OpenStudy (akashdeepdeb):

:D

OpenStudy (anonymous):

thankyou again